# Identifying efficient routes to laminarization: an optimization approach

### v1

@@ IV.2 Minimal seeds for relaminarization
Finally, in Fig. 11, the minimal seeds for relaminarization from multiple reference points in the turbulent region of the state space are shown. The reference points considered here are two points on the turbulent attractor – approximating the closest and farthest points on the attractor from the laminar state from a long, chaotic trajectory – and a point on an unstable periodic orbit in the turbulent region (shown in magenta in Fig. 11). This periodic orbit was found by Newton-Raphson iteration. These results show that the magnitude and direction of the optimal perturbation to trigger relaminarization vary depending on the chosen reference point in the turbulent region. However, each of the resulting trajectories follow a similar route to relaminarization, with all of them passing near the edge periodic orbit before laminarizing. Interestingly, two of these minimal seed trajectories initially move away from the laminar state, traveling around the attractor before laminarizing. These findings further underscore the counterintuitive nature of the optimal route to laminarization from the turbulent attractor, which is determined by a complex interaction with the boundary of the laminar basin of attraction.
@@ V Conclusion
In this work, we have proposed the concept of the minimal seed for relaminarization as a tool for analyzing the relaminarization behavior of chaotic or transiently chaotic fluid flows. We developed an optimization-based approach for approximating this critical perturbation of an attractor which leads to laminarization and demonstrated the effectiveness of this method on the nine-mode Galerkin model of a sinusoidal shear flow of Moehlis, et al. [25, 26]. This method yields a laminarizing trajectory that originates from an initial condition near the turbulent attractor and far from the laminar state in terms of energy. We expect that identifying such a trajectory will be useful for control, as it provides an reference pathway for achieving relaminarization. That is, if the minimal seed for relaminarization or a point on the resulting trajectory can be reached, the system will laminarize naturally without further actuation. Thus, the minimal seed for relaminarization offers a physically grounded and interpretable target for optimal drag-reduction control strategies.
@@
Additionally, in ongoing and future works, we aim to demonstrate the practical utility of these methods in the development of flow control strategies. A central focus will be to evaluate how control strategies developed using the minimal seed for relaminarization compare to established approaches for stabilizing turbulent flows. These include traditional methods, such as opposition control and more recent, advanced methods such as reinforcement learning. In addition to the development of new control strategies, the approach presented here can also be useful for the evaluation of existing control strategies, as the a comparison of the minimal seeds for actuated versus unactuated systems can serve as a measure of the robustness of a given control strategy.

### v2

@@ IV.2 Minimal seeds for relaminarization
In Fig. 11, the minimal seeds for relaminarization from multiple reference points in the turbulent region of the state space are shown. The reference points considered here are two points on the turbulent attractor – approximating the closest and farthest points on the attractor from the laminar state from a long, chaotic trajectory – and a point on an unstable periodic orbit in the turbulent region (shown in magenta in Fig. 11). This periodic orbit was found by Newton-Raphson iteration. These results show that although the optimal perturbation varies with the chosen reference point, each of the laminarizing trajectories are influenced by a common mechanism: they approach the edge periodic orbit and then depart along its unstable manifold toward the laminar state. Interestingly, two of these minimal seed trajectories initially move away from the laminar state, traveling around the attractor before laminarizing. These findings further underscore the counterintuitive nature of the optimal route to laminarization from the turbulent attractor, which is determined by a complex interaction with the boundary of the laminar basin of attraction.
IV.2.1 Application for control
We now illustrate how the minimal seed for relaminarization may be useful for developing a control sequence to drive the flow to the laminar state from a point on the chaotic attractor. The general idea is that the trajectory resulting from the minimal seed for relaminarization provides a helpful reference trajectory for controller design, as it originates near the turbulent attractor and laminarizes naturally without any actuation. Therefore, if we can actuate the system for a short time interval to steer it toward the laminarizing trajectory, then the system should continue to follow this trajectory to the laminar state without further actuation after the brief initial interval.
To illustrate this, we consider the reference point on the turbulent attractor from the results in Fig. 9-10. We consider a fully-actuated modification of the MFE system. That is, a control term is added to each of the states:
$\frac{da_{j}}{dt}=f_{j}(\mathbf{a})+v_{j}$ (17)
for $j=1,\dots,9$, where the $f_{j}(\mathbf{a})$ are the dynamics of the original MFE system and $v_{j}$ are the added control terms. With this system, we optimize a quadratic cost function, penalizing the deviation from the reference trajectory and the control effort as
$J=\int_{0}^{t_{F}}\left[(a-a_{\mathrm{ref}})^{\intercal}Q(a-a_{\mathrm{ref}})+v^{\intercal}Rv\right]dt+(a_{F}-(a_{\mathrm{ref}})_{F})^{\intercal}Q_{F}(a_{F}-(a_{\mathrm{ref}})_{F})$ (18)
where $t_{F}$ is the end of the actuation interval, $Q$, $R$, and $Q_{F}$ are scalar penalty weights on the in-horizon tracking error, the control, and the terminal state, $a_{F}=a(t_{F})$, $(a_{\mathrm{ref}})_{F}=a_{\mathrm{ref}}(t_{F})$, and other time-dependencies are dropped for brevity. For this demonstration, we use $Q=1$, $Q_{F}=10$, $R=1$, $t_{F}=50$, and set $a_{\mathrm{ref}}$ to be the resulting trajectory from the minimal seed for relaminarization computed above. To solve this problem numerically, we optimize the control inputs, $v_{j}$ at $50$ evenly spaced points between $t=0$ and $t=t_{F}$, holding the control constant between these points. The problem is solved numerically using torchdiffeq with an L-BFGS optimizer, as in the earlier optimization problems.
Figure 12: Control using the minimal seed for relaminarization. By applying control for a brief actuation interval of 50 time units to steer toward the minimal seed for relaminarization trajectory, the system is driven from a reference point, $\mathbf{x}_{T}$, on the turbulent attractor into the laminar basin of attraction. The controlled trajectory then laminarizes naturally after the actuation interval. A long chaotic trajectory is shown in gray to indicate the shape of the attractor.
Figure 13: Optimized control inputs to track the minimal seed trajectory used in Fig. 12.
Figure 12 shows the resulting solution from solving this optimal control problem. We see that by applying a control over a very short interval, the initial condition on the chaotic attractor is driven to follow the laminating trajectory from the minimal seed for relaminarization. After the short actuation interval, all of the control terms are set to zero and we see that the trajectory continues to laminarize naturally. Fig. 12 shows the uncontrolled trajectory on the attractor along with the controlled trajectory and the trajectory from the minimal seed for relaminarization, with a green marker on the controlled trajectory noting the point beyond which the actuation is turned off. Fig. 13 shows the optimized control inputs used to achieve this. This example demonstrates that the minimal seed for relaminarization and the resulting trajectory serve as a useful reference for solving an optimal control problem to laminarize the system, resulting in an efficient control strategy that only requires actuation for a short time to drive the system over the laminar-turbulent boundary.
@@ V Conclusion
In this work, we have proposed the concept of the minimal seed for relaminarization as a tool for analyzing the relaminarization behavior of chaotic or transiently chaotic fluid flows. We developed an optimization-based approach for approximating this critical perturbation of a reference point on an attractor which leads to laminarization and demonstrated the effectiveness of this method on the nine-mode Galerkin model of a sinusoidal shear flow of Moehlis, et al. [25, 26]. This method yields a laminarizing trajectory that originates from an initial condition near the turbulent attractor and far from the laminar state in terms of energy. We expect that identifying such a trajectory will be useful for control, as it provides an reference pathway for achieving relaminarization. That is, if the minimal seed for relaminarization or a point on the resulting trajectory can be reached, the system will laminarize naturally without further actuation. Thus, the minimal seed for relaminarization offers a physically grounded and interpretable target for optimal drag-reduction control strategies.
@@
Building on these ideas, we note that while the present study focuses solely on the reduced-order MFE system, our approach is conceptually general and provides a foundation for extensions to higher-dimensional models and direct numerical simulations of turbulent flows. Applying the method to full turbulence would require significant computational resources and additional modeling, but similar optimization frameworks have already been demonstrated in DNS studies of complex flows [37]. By establishing feasibility in a system that retains key transitional dynamics, this work serves as a proof of concept and highlights promising directions for future research, including investigations into more complex flow scenarios.
Additionally, in ongoing and future works, we aim to demonstrate the practical utility of these methods in the development of flow control strategies. A central focus will be to evaluate how control strategies developed using the minimal seed for relaminarization compare to established approaches for flow control. These include traditional methods, such as opposition control and more recent, advanced methods such as reinforcement learning. In addition to the development of new control strategies, the approach presented here can also be useful for the evaluation of existing control strategies, as the a comparison of the minimal seeds for actuated versus unactuated systems can serve as a measure of the robustness of a given control strategy.